OpenStudy (anonymous):

for each of the following forms determine whether the following limit type is indeterminate, ( I got 50% correct I am confused rest of them)

OpenStudy (anonymous):

limit type is indeterminate? As in an indeterminate form, or the limit doesn't exist?

OpenStudy (anonymous):

Please tell me which one i did wrong.

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

\[ \infty\cdot \infty = \infty \]

OpenStudy (anonymous):

\(\infty^1=\infty\)

OpenStudy (anonymous):

so its IND right

OpenStudy (anonymous):

\[ 0/\infty=0 \]

OpenStudy (anonymous):

IND means infinite or indeterminate? Because they are not the same!

OpenStudy (anonymous):

indeterminate. I posted the question pic above. please look

OpenStudy (anonymous):

\[ \pi^{-\infty}=1/\pi^{\infty} = 1/\infty = 0 \]

OpenStudy (anonymous):

Then I guess they want you to put DNE if it goes to infinity.

OpenStudy (anonymous):

I fixed it still 50% :/

OpenStudy (anonymous):

its

OpenStudy (anonymous):

Mention some other people for help.

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

ok, Thank you

OpenStudy (abb0t):

\(\frac{ ∞ }{ 0 } = \) IND

OpenStudy (anonymous):

There are \(7\) indeterminate forms: \[ \frac{0}{0}, 0\cdot \infty, \frac{\infty}{\infty}, \infty - \infty, 0^0, \infty^0, 1^\infty \]

OpenStudy (anonymous):

Kinda busy filling up my notebook now. Some of the data was lost before.

OpenStudy (anonymous):

ok i will check my answer if i made some mistake

OpenStudy (anonymous):

thanks wio!

OpenStudy (anonymous):

@electrokid